one type of punch is 20% ginger ale. and another type of punch is 60% ginger ale. How much of each kind should be used to make 100 gallons of punch that is 30% ginger
Answer by TimothyLamb(4379)

---
20x + 60y = 30*100
x + y = 100
---
20x + 60y = 3000
x + y = 100
---
copy and paste the above standard form linear equations in to this solver:
https://sooeet.com/math/system-of-linear-equations-solver.php
---
answer:
x= 75 gal of 20%
y= 25 gal of 60%
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php